An overworked teacher placed ethanol and ethyl bromide in two different bottles but forget to label them.What simple chemical test is performed on a sample from each bottle to identify them???
1
Expert's answer
2013-02-28T09:20:58-0500
The simplest chemical test for it is reaction with NaOH

Ethanol would not show some sign of reaction.

Ethyl bromide would react with producing white precipitate of NaBr which is not soluble in CH3-CH2-Br.

CH3-CH2-Br + NaOH = NaBr + CH2=CH2 + H2O